# Identifying bimodality in data using a proximity-based null model with an application to classifying cell cycle phases using oxidative stress

### v2

@@ V Identifying cell states using oxidative stress
Inspired by the possibility of identifying addition sub-phases of the data, we close by examining a new data set taken from the same group of cells, but 3.5 hours later on, i.e., with DFC and CF measurements taken at 5.5. hours after release. At this point it is likely that all cells are no longer mitotic and are in some sub-phase of interphase, specifically either the S or G2 phase. (Note that S and G2 correspond to the phases where, respectively, the cells are either replicating DNA and synthesizing protein in preparation for the next mitosis phase. Moreover, Dictyostelium discoideum does not pass through a G1 phase.) Importantly, S phase is much shorter than G2 phase Weeks1994 , and therefore is likely to correspond to a smaller group of cells at any given time. Furthermore, investigations into cancer cells show that DCF/CF ratios measuring oxidative stress tend to be smaller in S phase as compared to G2 phase Patterson2019 .
@@ Appendix A Experimental measurements and data collection
The data set considered above in Sec. V consists of Dictyostelium discoideum cells for which we consider the measurements of oxidative stress. To quantify oxidative stress, we compare the levels of fluorescence from dihydrodichlorofluorescein diacetate (DCF) to carboxyfluorescein diacetate (CF) after initially loading each cell with $250$ $\mu$M of each. In particular, CF acts as an internal standard while DCF is a reactive oxygen species probe, indicating oxidative stress. Additionally, each cell is also initially loaded with 40 $\mu$M Rose Bengal as a photosensitizer to induce oxidative stress. DCF is a common indicator for reactive oxygen species Chen2010 . Recent research suggests that ratioed measurements with CF more accurately reflect cell-to-cell variation in reactive oxygen species by accounting for spurious sources of heterogeneity Metto2013 ; Rodogiannis2018 ; Sibbitts2020 ; Allcroft2023
We note that heterogeneities in the individual measurements of CF and DCF are significant, so overall oxidative stress is best gauged using their log-ratio, yielding a single value for each cell. After incubation and washing in low-fluorescence media, cell are exposed to a blue LED light source for a time ranging between 0-10 minutes, after which the log ratio $x=\text{log}(\text{DCF}/\text{CF})$ is measured as an individual measurement of oxidative stress. Additional details can be found in Ref. Allcroft2023 .
Acknowledgements.
MLK and PSS acknowledge support from NSF grant MCB-2126177.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.

### v3

@@ 5 Identifying cell states using oxidative stress
Inspired by the possibility of identifying addition sub-phases of the data, we close by examining a new data set taken from the same group of cells, but 3.5 hours later on, i.e., with DFC and CF measurements taken at 5.5. hours after release. At this point it is likely that all cells are no longer mitotic and are in some sub-phase of interphase, specifically either the S or G2 phase. (Note that S and G2 correspond to the phases where, respectively, the cells are either replicating DNA and synthesizing protein in preparation for the next mitosis phase. Moreover, Dictyostelium discoideum does not pass through a G1 phase.) Importantly, S phase is much shorter than G2 phase [29], and therefore is likely to correspond to a smaller group of cells at any given time. Furthermore, investigations into cancer cells show that DCF/CF ratios measuring oxidative stress tend to be smaller in S phase as compared to G2 phase [30].
@@ 6 Discussion
[…]
Author contributions MLK, TJA, and PSS performed all work associated with this manuscript.
[…]
Funding This work was funded by NSF grant MCB-2126177 (MLK and PSS).
[…]
Availability of Data and Materials Data and materials are available on reasonable request from PSS.
Declarations
Competing Interests
The authors declare that they have no conflict of interest.
@@ Appendix A Experimental measurements and data collection
The data set considered above in Sec. 5 consists of Dictyostelium discoideum cells for which we consider the measurements of oxidative stress. To quantify oxidative stress, we compare the levels of fluorescence from dihydrodichlorofluorescein diacetate (DCF) to carboxyfluorescein diacetate (CF) after initially loading each cell with $250$ $\mu$M of each. In particular, CF acts as an internal standard while DCF is a reactive oxygen species probe, indicating oxidative stress. Additionally, each cell is also initially loaded with 40 $\mu$M Rose Bengal as a photosensitizer to induce oxidative stress. DCF is a common indicator for reactive oxygen species [31]. Recent research suggests that ratioed measurements with CF more accurately reflect cell-to-cell variation in reactive oxygen species by accounting for spurious sources of heterogeneity [32, 33, 34, 19]
We note that heterogeneities in the individual measurements of CF and DCF are significant, so overall oxidative stress is best gauged using their log-ratio, yielding a single value for each cell. After incubation and washing in low-fluorescence media, cell are exposed to a blue LED light source for a time ranging between 0-10 minutes, after which the log ratio $x=\text{log}(\text{DCF}/\text{CF})$ is measured as an individual measurement of oxidative stress. Additional details can be found in Ref. [19].